OpenStudy (anonymous):

What is the vertex of the graph of y=x^2-11x+28

OpenStudy (anonymous):

vertex x=-b/2a x=-(-11)/2(1)=11/2=5.5 ans

OpenStudy (anonymous):

subs this x=5.5 to y=x^2-11x+28 we get y=(5.5)^2-11(5.5)+28 y=-9/4=-2.25 ans vertex=(x,y)=(11/2,-9/4) ans